OpenStudy (i_love_my_nieces):

1+1 Real question in comments

OpenStudy (i_love_my_nieces):

I have the answer I just don't know how to work it out.\[\frac{ 1 }{ 2 }(b-3)=\frac{ 5 }{ 2}\]

OpenStudy (i_love_my_nieces):

\[b = \frac{ 4 }{ 0.5 } \]

OpenStudy (anonymous):

o.o..

OpenStudy (i_love_my_nieces):

But\[\frac{ 4 }{ 0.5 }=8\] in simplest form

OpenStudy (i_love_my_nieces):

@Teddyiswatshecallsme

Nnesha (nnesha):

whats the answer ?

OpenStudy (i_love_my_nieces):

@Nnesha 8

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @I_love_my_nieces I have the answer I just don't know how to work it out.\[\frac{ 1 }{ 2 }(b-3)=\frac{ 5 }{ 2}\] \(\color{blue}{\text{End of Quote}}\) how would you get rid of the 2 at left side would you add , subtract multiply or divide ?

OpenStudy (i_love_my_nieces):

The 2 at what end?

OpenStudy (i_love_my_nieces):

@Nnesha

Nnesha (nnesha):

multiply 2 both sides to get rid of 2

OpenStudy (i_love_my_nieces):

Where do you see a 2?

Nnesha (nnesha):

\[\frac{ 1 }{ \color{ReD}{2} }(b-3)=\frac{ 5 }{ 2}\]

Nnesha (nnesha):

that's 2 ^^

Nnesha (nnesha):

we have to solve for b right ? so we need to move all the numbers which is at left side to the right side

Nnesha (nnesha):

so that's why multiply both sides by 2 to cancel out the red two at left side \[\frac{ 1 }{ \color{ReD}{2} }(b-3)=\frac{ 5 }{ 2}\]

OpenStudy (i_love_my_nieces):

SO multiply 1/2 by 2? and 5/2 by 2?

Nnesha (nnesha):

yes right!

OpenStudy (i_love_my_nieces):

1 and 5

Nnesha (nnesha):

yes right so \[\huge\rm 1(b-3)=5 ~~~~~~~~b-3=5\] now can you solve for b ?

OpenStudy (i_love_my_nieces):

brb stay right here

Nnesha (nnesha):

o_0 i hve 2 mints XD

Nnesha (nnesha):

knock knock i'm leaving ..... :D

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha yes right so \[\huge\rm 1(b-3)=5 ~~~~~~~~b-3=5\] now can you solve for b ? \(\color{blue}{\text{End of Quote}}\) after this move the -3 to the right side sorry i have to go

Nnesha (nnesha):

if you have question about this please ask Luigi0210 :D

OpenStudy (i_love_my_nieces):

@Luigi0210

Nnesha (nnesha):

got it ?